# Supersonic–sonic patch solution for the two-dimensional Euler equations with a van der Waals equation of state 11footnotemark: 1

### v1

@@ 4.1 Inversion
$\partial_{-}\xi=\frac{\kappa(\tau)t(\cos\theta t+\sin\theta\sqrt{1-t^{2}})}{\sqrt{1-t^{2}}f},\;\partial_{-}\eta=\frac{\kappa(\tau)t(\sin\theta t+\cos\theta\sqrt{1-t^{2}})}{\sqrt{1-t^{2}}g}.$ (4.3)
@@
$\begin{split}\xi(z^{0},t^{0})&=\xi^{*}(\tilde{z}(t^{*}))-\int_{t^{0}}^{t^{*}}\frac{\kappa(\tau)t(\cos\theta t+\sin\theta\sqrt{1-t^{2}})}{\sqrt{1-t^{2}}f}(z_{-}(t;t^{0},z^{0}),t)dt,\\ \eta(z^{0},t^{0})&=\psi(\xi^{*}(\tilde{z}(t^{*})))-\int_{t^{0}}^{t^{*}}\frac{\kappa(\tau)t(\sin\theta t+\cos\theta\sqrt{1-t^{2}})}{\sqrt{1-t^{2}}g}(z_{-}(t;t^{0},z^{0}),t)dt.\end{split}$
@@
For an arbitrary point $(z^{0},t^{0})$ in the region $M^{\prime}N^{\prime}O^{\prime}$, we trace the negative characteristic curve $z=z_{-}(t;t^{0},z^{0})$ for $t\geq t^{0}$ until it intersects the boundary $M^{\prime}O^{\prime}$ at a unique point $(\tilde{z}(t^{\ast}),t^{\ast})$ satisfying
$\begin{split}\begin{cases}\frac{dz_{-}}{dt}(t;t^{0},z^{0})&=-\frac{cft^{2}}{Y-tg}(z_{-}(t;t^{0},z^{0}),t),\\ z_{-}(t^{0};t^{0},z^{0})&=z^{0},\\ z_{-}(t^{*};t^{0},z^{0})&=\tilde{z}(t^{*}).\end{cases}\end{split}$
We integrate (4.4) along the negative characteristic $z=z_{-}(t;t^{0},z^{0})$, starting from the point $(z^{0},t^{0})$ and proceeding to the boundary point $(\tilde{z}(t^{\ast}),t^{\ast})$. By incorporating the prescribed boundary values of $\theta$ on $\widehat{M^{\prime}O^{\prime}}$, we obtain
$\theta(z^{0},t^{0})=\theta^{*}(\xi^{*}(\tilde{z}(t^{*})))+\int_{t^{0}}^{t^{*}}\frac{ft^{2}\sqrt{1-t^{2}}}{Y-gt}\left(-\kappa(\tau)t\sqrt{1-t^{2}}Y+\frac{\bar{\omega}^{2}}{c}\right)(z_{-}(t;t^{0},z^{0}),t)dt.$
@@ 4.2 Mapping $(z,t)\mapsto(\xi,\eta)$ is one to one
$(\phi_{\xi},\phi_{\eta})\cdot(\bar{\omega}_{\eta},-\bar{\omega}_{\xi})=-\frac{c({X}-{Y})}{2t}(1+\kappa(\tau)(1-t^{2}))<0.$ (4.5)
@@
where the region $MNO$ is enclosed by the curves $\widehat{MN}$, $\widehat{NO}$ and $\widehat{MO}$. In particular, the curve $\widehat{MN}$ is characterized by
$\widehat{MN}=\{(\xi,\eta)\;|\;\bar{\omega}(\xi,\eta)=1,\;\xi\in[\xi^{*},\xi_{2}]\},$
where $\xi^{\ast}=\xi(\bar{z}(0),0)$ and the curve $\widehat{NO}$ is specified by
$\widehat{NO}=\{(\xi,\eta)\;|\;\hat{z}(\xi,\eta)=z_{+}(\hat{t}(\xi,\eta);\bar{\delta},\tilde{z}(\bar{\delta})),\;\xi\in[\xi^{**},\xi^{*}]\},$ (4.8)
where the function $z_{+}(t;\bar{\delta},\tilde{z}(\bar{\delta}))$ denotes the solution to the following ordinary differential equation
@@
Also, the number $\xi^{**}$ satisfies
$\tilde{z}(\xi^{**},\psi(\xi^{**}))=z_{+}(\tilde{t}(\xi^{**},\psi(\xi^{**}));\bar{\delta},\tilde{z}(\bar{\delta})).$
In the $(\xi,\eta)$ plane, the points $N$ and $O$ have coordinates $(\xi^{\ast},\eta(\bar{z}(0),0))$ and $(\xi^{\ast\ast},\psi(\xi^{\ast\ast}))$, respectively. By the construction of the coordinate functions $(\xi(z,t),\eta(z,t))$, it follows that the functions $(c(\xi,\eta),\theta(\xi,\eta),\bar{\omega}(\xi,\eta))$ defined in (4.7) satisfy the boundary conditions prescribed in (2.22). Based on (4.7), we further introduce the functions $(\omega,\alpha,\beta)(\xi,\eta)$ for all $(\xi,\eta)\in\overline{MNO}$ as follows
$\begin{split}\omega(\xi,\eta)=\arcsin\bar{\omega}(\xi,\eta),\;\alpha(\xi,\eta)=\theta(\xi,\eta)+\omega(\xi,\eta),\;\beta(\xi,\eta)=\theta(\xi,\eta)-\omega(\xi,\eta).\end{split}$ (4.10)
@@
In view of (4.11) and (4.2), it is sufficient to establish that the functions $(X,Y,W)(\xi,\eta)$ are uniformly $C^{\mu}$- continuous for any $\mu\in(0,1/3)$ throughout the region $MNO$. These regularity properties follow from Lemma 3.6 together with the following assertion. If a function $I(z,t)$ belongs to $C^{2,\tilde{\mu}}$ on the entire region $M^{\prime}N^{\prime}O^{\prime}$, then the transformed function $\tilde{I}(\xi,\eta):=I(\hat{z}(\xi,\eta),\hat{t}(\xi,\eta))$ is uniformly $C^{\tilde{\mu}}$ continuous on $MNO$.
To justify this claim, let $(\xi^{\prime},\eta^{\prime})$ and $(\xi^{\prime\prime},\eta^{\prime\prime})$ be two arbitrary points in $MNO$ and let $(z^{\prime},t^{\prime})$ and $(z^{\prime\prime},t^{\prime\prime})$ denote their corresponding points in the $(z,t)$ plane. We then examine the differences
$|z^{\prime}-z^{\prime\prime}|=|\phi(\xi^{\prime},\eta^{\prime})-\phi(\xi^{\prime\prime},\eta^{\prime\prime})|\leq M_{1}(|\xi^{\prime\prime}-\xi^{\prime}|+|\eta^{\prime\prime}-\eta^{\prime}|),$ (4.12)
@@
$\begin{split}|t^{\prime}-t^{\prime\prime}|^{2}&\leq|t^{\prime}-t^{\prime\prime}||t^{\prime}+t^{\prime\prime}|=|t^{\prime 2}-t^{\prime\prime 2}|=|\bar{\omega}^{2}(\xi^{\prime},\eta^{\prime})-\bar{\omega}^{2}(\xi^{\prime\prime},\eta^{\prime\prime})|\\ &\leq 2|\bar{\omega}(\xi^{\prime},\eta^{\prime})-\bar{\omega}(\xi^{\prime\prime},\eta^{\prime\prime})|\leq M_{2}(|\xi^{\prime\prime}-\xi^{\prime}|+|\eta^{\prime\prime}-\eta^{\prime}|),\end{split}$ (4.13)
@@
$M_{1}=\max_{(\xi,\eta)\in MNO}\frac{c}{\bar{\omega}}(\xi,\eta),\;M_{2}=2\max\left\{\max_{(\xi,\eta)\in MNO}|\bar{\omega}_{\xi}|,\;\max_{(\xi,\eta)\in MNO}|\bar{\omega}_{\eta}|\right\}.$
Here $M_{1}$ and $M_{2}$ denote two uniform positive constants determined by (4.2) and Lemmas 3.1 and 3.4. Combining these bounds with (4.12) and (4.13) and invoking the $C^{2\tilde{\mu}}$ continuity, we obtain
[…]
for some uniform positive constant $M_{3}>0$, which implies that the function $\tilde{I}(\xi,\eta)$ is uniformly $C^{\tilde{\mu}}$ continuous throughout the region $MNO$. By Lemma 3.6, we further conclude that $(X,Y)(\xi,\eta)$ are uniformly $C^{(1-\nu)/2}$ continuous and that $W(\xi,\eta)$ is uniformly $C^{(2-\nu)/6}$ continuous on $MNO$ for any $\nu\in(0,1)$. Consequently, $(X,Y,W)(\xi,\eta)$ is uniformly $C^{\mu}$ continuous for all $\mu\in(0,1/3)$.
@@
We differentiate $\hat{z}(\xi,\eta)=z^{+}(\hat{t}(\xi,\eta);\bar{\delta},\tilde{z}(\bar{\delta}))$ with respect to $\xi$ and apply (4.9), we obtain
@@
Combining Lemmas 4.1-4.3, we complete the proof of Theorem 2.2. Using (4.7) and (2.8), we now define the functions $(c,u,v)(\xi,\eta)$ as follows:

### v2

@@ 4.1 Inversion
$\partial_{-}\xi=-\frac{t^{2}f(\sqrt{1-t^{2}}\sin\theta+ct\cos\theta)}{Y-tg},\;\partial_{-}\eta=\frac{t^{2}f(\sqrt{1-t^{2}}\cos\theta-ct\sin\theta)}{Y-tg}.$ (4.3)
@@
$\begin{split}\xi(z^{0},t^{0})&=\xi^{*}(\tilde{z}(t^{*}))+\int_{t^{0}}^{t^{*}}\frac{t^{2}f(\sqrt{1-t^{2}}\sin\theta+ct\cos\theta)}{Y-tg}(z_{-}(t;t^{0},z^{0}),t)dt,\\ \eta(z^{0},t^{0})&=\psi(\xi^{*}(\tilde{z}(t^{*})))-\int_{t^{0}}^{t^{*}}\frac{t^{2}f(\sqrt{1-t^{2}}\cos\theta-ct\sin\theta)}{Y-tg}(z_{-}(t;t^{0},z^{0}),t)dt.\end{split}$
@@
For an arbitrary point $(z_{0},t_{0})$ in the region $M^{\prime}N^{\prime}O^{\prime}$, we trace the negative characteristic curve $z=z_{-}(t;t_{0},z_{0})$ for $t\geq t_{0}$ until it intersects the boundary $M^{\prime}O^{\prime}$ at a unique point $(\tilde{z}(t^{\ast}),t^{\ast})$ satisfying
$\begin{split}\begin{cases}\frac{dz_{-}}{dt}(t;t_{0},z_{0})&=-\frac{cft^{2}}{Y-tg}(z_{-}(t;t_{0},z_{0}),t),\\ z_{-}(t_{0};t_{0},z_{0})&=z_{0},\\ z_{-}(t^{*};t_{0},z_{0})&=\tilde{z}(t^{*}).\end{cases}\end{split}$
We integrate (4.4) along the negative characteristic $z=z_{-}(t;t_{0},z_{0})$, starting from the point $(z_{0},t_{0})$ and proceeding to the boundary point $(\tilde{z}(t^{\ast}),t^{\ast})$. By incorporating the prescribed boundary values of $\theta$ on $\widehat{M^{\prime}O^{\prime}}$, we obtain
$\theta(z_{0},t_{0})=\theta^{*}(\xi^{*}(\tilde{z}(t^{*})))+\int_{t_{0}}^{t^{*}}\frac{ft^{2}\sqrt{1-t^{2}}}{Y-gt}\left(-\kappa(\tau)t\sqrt{1-t^{2}}Y+\frac{\bar{\omega}^{2}}{c}\right)(z_{-}(t;t_{0},z_{0}),t)dt.$
@@ 4.2 Mapping $(z,t)\mapsto(\xi,\eta)$ is one to one
$\nabla\phi\cdot(\bar{\omega}_{\eta},-\bar{\omega}_{\xi})=-\frac{c({X}-{Y})}{2t}(1+\kappa(\tau)(1-t^{2}))<0.$ (4.5)
@@
where the region $MNO$ is enclosed by the curves $\widehat{MN}$, $\widehat{NO}$ and $\widehat{MO}$. In particular, the curves $\widehat{MN}$ and $\widehat{NO}$ are characterized by
$\begin{split}\widehat{MN}&=\{(\xi,\eta)\;|\;\bar{\omega}(\xi,\eta)=1,\;\xi\in[\xi^{\prime},\xi_{2}]\},\\ \widehat{NO}&=\{(\xi,\eta)\;|\;\hat{z}(\xi,\eta)=z_{+}(\hat{t}(\xi,\eta);\bar{\delta},\tilde{z}(\bar{\delta})),\;\xi\in[\xi^{\prime\prime},\xi^{\prime}]\},\end{split}$ (4.8)
where the point $\xi^{\prime}=\xi(\bar{z}(0),0)$ and the function $z_{+}(t;\bar{\delta},\tilde{z}(\bar{\delta}))$ denotes the solution to the following ordinary differential equation
@@
Also, $\xi^{\prime\prime}$ satisfies
$\hat{z}(\xi^{\prime\prime},\psi(\xi^{\prime\prime}))=z_{+}(\tilde{t}(\xi^{\prime\prime},\psi(\xi^{\prime\prime}));\bar{\delta},\tilde{z}(\bar{\delta})).$
In the self-similar plane, the points $N$ and $O$ have coordinates $(\xi^{\prime},\eta(\bar{z}(0),0))$ and $(\xi^{\prime\prime},\psi(\xi^{\prime\prime}))$, respectively. By the construction of the coordinate functions $(\xi(z,t),\eta(z,t))$, it follows that the functions $(c(\xi,\eta),\theta(\xi,\eta),\bar{\omega}(\xi,\eta))$ defined in (4.7) satisfy the boundary conditions prescribed in (2.22). Based on (4.7), we further introduce the functions $(\alpha,\omega,\beta)(\xi,\eta)$ for all $(\xi,\eta)\in\overline{MNO}$ as follows
$\alpha(\xi,\eta)=\theta(\xi,\eta)+\omega(\xi,\eta),\;\omega(\xi,\eta)=\arcsin\bar{\omega}(\xi,\eta),\;\beta(\xi,\eta)=\theta(\xi,\eta)-\omega(\xi,\eta).$ (4.10)
@@
In view of (4.11) and (4.2), it is sufficient to establish that the functions $(X,Y,W)(\xi,\eta)$ are uniformly $C^{\mu}$- continuous for any $\mu\in(0,1/3)$ throughout the region $MNO$. These regularity properties follow from Lemma 3.6 together with the following assertion. If a function $I(z,t)$ belongs to $C^{2\bar{\mu}}$ on the entire region $M^{\prime}N^{\prime}O^{\prime}$, then the transformed function $\bar{I}(\xi,\eta):=I(\hat{z}(\xi,\eta),\hat{t}(\xi,\eta))$ is uniformly $C^{\bar{\mu}}$ continuous on $MNO$.
To justify this claim, let $(\xi^{*},\eta^{*})$ and $(\xi^{**},\eta^{**})$ be two arbitrary points in $MNO$ and let $(z^{*},t^{*})$ and $(z^{**},t^{**})$ denote their corresponding points in the $(z,t)$ plane. We then examine the differences
$|z^{*}-z^{**}|=|\phi(\xi^{*},\eta^{*})-\phi(\xi^{**},\eta^{**})|\leq a_{7}(|\xi^{**}-\xi^{*}|+|\eta^{**}-\eta^{*}|),$ (4.12)
@@
$\begin{split}|t^{*}-t^{**}|^{2}&\leq|t^{*}-t^{**}||t^{*}+t^{**}|=|{t^{*}}^{2}-{t^{**}}^{2}|=|\bar{\omega}^{2}(\xi^{*},\eta^{*})-\bar{\omega}^{2}(\xi^{**},\eta^{**})|\\ &\leq 2|\bar{\omega}(\xi^{*},\eta^{*})-\bar{\omega}(\xi^{**},\eta^{**})|\leq a_{8}(|\xi^{**}-\xi^{*}|+|\eta^{**}-\eta^{*}|),\end{split}$ (4.13)
@@
$a_{7}=\max_{(\xi,\eta)\in MNO}\frac{c}{\bar{\omega}}(\xi,\eta),\;a_{8}=2\max\left\{\max_{(\xi,\eta)\in MNO}|\bar{\omega}_{\xi}|,\;\max_{(\xi,\eta)\in MNO}|\bar{\omega}_{\eta}|\right\}.$
Here $a_{7}$ and $a_{8}$ denote two positive constants determined by (4.2) and Lemmas 3.1 and 3.4. Combining these bounds with (4.12) and (4.13) and invoking the $C^{2\bar{\mu}}$ continuity, we obtain
$\begin{split}|\tilde{I}(\xi^{*},\eta^{*})-\tilde{I}(\xi^{**},\eta^{**})|&=|I(z^{*},t^{*})-I(z^{**},t^{**})|\\ &\leq a_{9}|(z^{*},t^{*})-(z^{**},t^{**})|^{2\bar{\mu}}\\ &\leq a_{9}\left(|z^{*}-z^{**}|^{2}+|t^{*}-t^{**}|^{2}\right)^{\bar{\mu}}\\ &\leq a_{9}|(\xi^{*},\eta^{*})-(\xi^{**},\eta^{**})|^{\bar{\mu}},\end{split}$
for some uniform positive constant $a_{9}$, which implies that the function $\tilde{I}(\xi,\eta)$ is uniformly $C^{\bar{\mu}}$ continuous throughout the region $MNO$. By Lemma 3.6, we further conclude that $(X(\xi,\eta),Y(\xi,\eta))$ are uniformly $C^{(1-\nu)/2}$ continuous and that $W(\xi,\eta)$ is uniformly $C^{(2-\nu)/6}$ continuous on $MNO$ for any $\nu\in(0,1)$. Consequently, $(X(\xi,\eta),Y(\xi,\eta),W(\xi,\eta))$ is uniformly $C^{\mu}$ continuous for all $\mu\in(0,1/3)$.
@@
Let $\hat{z}(\xi,\eta)=z^{+}(\hat{t}(\xi,\eta);\bar{\delta},\tilde{z}(\bar{\delta}))$. Differentiating it with respect to its first argument and applying (4.9), we obtain
@@
Combining Lemmas 4.1-4.3, we complete proof of the Theorem 2.2. Using (4.7) and (2.8), we now define functions $(c,u,v)(\xi,\eta)$ as follows: